A. 10kW transmitter emits radio waves of 500 m wavelength .The no. of photon emitted per second by transmitter is of the order

Grade:12

1 Answers

Sanju
6 years ago
Energy emitted by the transmitter per sec,Et = 10kW
Energy of the wave, Ew = hc/λ 
where λ= 500m
Plugging the values gives Ew = 3.96 * 10-28
No of photons emitted per sec, n = Et / E= 10000 / 3.96 * 10-28 = 2.5 * 1031